20 men can reap a field in 20 day.when should 5 men leave the work,if the whole field is to be reapedgin 24 day after they leave the work.
Answer by ankor@dixie-net.com(22740)

20 men can reap a field in 20 days.
Find how many man-days to do the job.
20 * 20 = 400 man-days
:
when should 5 men leave the work, if the whole field is to be reaped in 24 days after they leave the work.
let d = no. of days until the remaining 15 men finish in 24 days
5d + 15(d+24) = 400
5d + 15d + 360 = 400
20d = 400 - 360
20d = 400 - 360
20d = 40
d = 40/20
d = 2 days
:
The 5 men can leave in 2 days
:
:
See if that checks out
5 * 2 = 10 man-days
15(2 + 24) = 390 man-days
---------------------
total: 400 man-days